not available, but the following data were calculated from a sample of 10 ob- servations: 0 Σ XY= 2,246 y? 41,208 Σ ? = 310 Σ xy=-654 y,2 ,-7,568 Σ.x2 = 60 (a) What is by? (b) What is bo?
Answer #1

Answer:

a) bi = (10*(-654)-(0*0))/ (10*3 10-02) = 0.1 *(0-0) =0